# Memorization to Generalization: Emergence of Diffusion Models from Associative Memory

### v2

@@ 2 Diffusion Models and Dense Associative Memories
where we also defined the energy $E^{\mathrm{DM}}$ of diffusion model, which up to state– or ${\mathbf{x}}$–independent terms is equal to
$E^{\text{DM}}({\mathbf{x}}_{t},t)=-2\sigma^{2}t\log[\sum\limits_{\mu=1}^{K}\exp\Big(-\frac{\lVert{\mathbf{x}}_{t}-\bm{\xi}^{\mu}\rVert^{2}_{2}}{2\sigma^{2}t}\Big{missing})\bigg{]}$ (5)
As already observed in [37], this energy function (5) is closely related to the class of models called DenseAMs, or Modern Hopfield Networks, which are large memory storage variants of classical Hopfield networks [4, 62].
The core idea behind DenseAMs is to design an energy function that sharply peaks around the intended memory patterns. These networks can store and retrieve a much larger number of memories, compared to the classical Hopfield networks, which scales super-linearly (and possibly exponentially) to the size of the network and decouples information storage capacity from the dimensionality of the data [4, 6]. Of particular interest here is the DenseAM model studied in [12] (see also [11]):
$E^{\text{AM}}({\mathbf{x}})=-\beta^{-1}\log[\sum\limits_{\mu=1}^{K}\exp\Big(-\beta\lVert{\mathbf{x}}-\bm{\xi}^{\mu}\rVert^{2}_{2}\Big{missing})\bigg{]}$ (6)
where $\beta$ is the inverse “temperature”, which controls the steepness of the energy landscape around memories $\bm{\xi}^{\mu}$. A direct comparison of the energy formulas (5) and (6) highlights a close connection between these two frameworks. Specifically, the data points in the diffusion model framework play the role of memories in the associative memory formulation. The variance of the noise added during the forward diffusion process (1) plays the role of the effective temperature $\beta^{-1}$. Finally, the reverse denoising process (2) corresponds to memory retrieval dynamics [36, 37].
3 Memorization-to-Generalization in 2D
@@
To better illustrate this connection, it is instructive to investigate a simple 2-dimensional toy model, see Figure 3, which exhibits many aspects of the memorization-generalization transition. Imagine that the training data lies on a unit circle. We are interested in exploring how the shape of the energy function (6) changes as the number of training data points increases.
$\bm{[K=1]}$ For a single data point $\bm{\xi}^{1}$, the energy function (6) has a single local minimum at that stored pattern.
$\begin{split}E^{\text{AM}}({\mathbf{x}})&=-\beta^{-1}\log[\exp\Big(-\beta\lVert{\mathbf{x}}-\bm{\xi}^{1}\rVert^{2}_{2}\Big{missing})\Big{]}=\lVert{\mathbf{x}}-\bm{\xi}^{1}\rVert^{2}_{2}\end{split}$ (7)
The shape of the energy landscape is independent of the value of $\beta$.
Figure 3: Energy landscape evolution for the 2D toy model as training data size $K$ increases. Models trained at $K\in\{2,9,1000\}$ use standard VE-SDE based diffusion pipeline with training data sampled from the unit circle, shown in black for $K\in\{2,9\}$. Generated samples are shown alongside the learned score field $s_{\theta}({\mathbf{x}}_{t},t)$ done via a neural network, aligned with the negative gradient of the energy (5). Hierarchical clustering identifies structure within the generations, with cluster centroid energies visualized by color and numerical value. The rightmost panel shows the exact solution as $K\rightarrow\infty$ derived in Equation 12. As $K$ grows, the model initially memorizes individual data points, forming isolated basins. Around $K=9$, spurious patterns emerge — distinct low-energy attractors not present in the data — which mark the onset of generalization. At large $K$, the model enters a fully generalized regime, where low-energy states lie on a flat, continuous manifold shown in Figure 1.
$\bm{[K=2]}$ For two patterns, the energy is given by
$\begin{split}E^{\text{AM}}({\mathbf{x}})=-\beta^{-1}&\log[\exp\Big(-\beta\lVert{\mathbf{x}}-\bm{\xi}^{1}\rVert^{2}_{2}\Big{missing})+\exp\Big(-\beta\lVert{\mathbf{x}}-\bm{\xi}^{2}\rVert^{2}_{2}\Big{missing})\Big{]}\end{split}$ (8)
For $\beta\rightarrow\infty$, this energy has two local minima corresponding to the stored patterns. Meanwhile, for finite values of $\beta$, there exist configurations of patterns such that the energy has only one local minimum:
$\bm{\eta}=\operatorname*{arg\,min}_{{\mathbf{x}}}E^{\text{AM}}({\mathbf{x}})$ (9)
such that this local minimum is different from either of the two stored patterns: $\bm{\eta}\neq\bm{\xi}^{1}$ and $\bm{\eta}\neq\bm{\xi}^{2}$. This “emergent” local minimum is the spurious state, see also the cartoon illustration in Figure 1.
@@
$\bm{[K\rightarrow\infty]}$ Finally, consider the case when the number of training data points is infinite. In this case they can be described by a continuous density of states:
$p({\mathbf{y}})=\frac{1}{\pi}\delta\big{(}y_{1}^{2}+y_{2}^{2}-1\big{)}$ (10)
@@
[…]
@@
where $I_{0}(\cdot)$ is a modified Bessel function of the first kind, see Appendix A for more details on the derivation of Equation 11. Thus, the energy of the model is given by
@@
$\begin{split}E^{\text{AM}}(R,\phi)&=R^{2}+1-\frac{1}{\beta}\log[I_{0}(2\beta R)\big{]}\underset{\beta\rightarrow\infty}{\approx}(R-1)^{2}\end{split}$ (12)
where $R$ is the radius of the unit circle and $\phi$ is the polar angle.

### v3

@@ 3  Discussion
4  Results
The main result of our work is the empirical demonstration that DMs trained with standard methods generate spurious samples at the onset of memorization-to-generalization transition. The existence of these samples is a cornerstone of energy-based AM models. By casting DMs’ generation step into DenseAM setting, one can theoretically expect the emergence of these states at the transition boundary. We have characterized computational properties of memorized, spurious, and generalized samples by studying their abundance, volume of the basins of attraction, and curvature of the energy landscape as a function of the DM training set size. Spurious samples in DMs have well-defined basins of attraction around them and have not been discussed in prior literature.
Our theoretical arguments pertain to the idealistic setting in which the neural network $s_{\theta}({\mathbf{x}}_{t},t)$ is capable of learning the true empirical distribution obtained by adding noise to the training data. This limiting case is often adopted in theoretical studies of DMs, see Biroli et al. (2024); Biroli and Mézard (2023); Achilli et al. (2025); Vastola (2025). From a practical perspective, this is achieved if $s_{\theta}({\mathbf{x}}_{t},t)$ is over-parametrized and trained for a long time. This simplification allows us to make a theoretical prediction that DMs must generate spurious states. In practical settings, the size of the score modeling neural network is finite. It is non-trivial that the theoretical prediction obtained in the regime of an infinitely large neural network can be extrapolated to the regime in which the size of $s_{\theta}({\mathbf{x}}_{t},t)$ is finite and the amount of training data is large. A possible perspective on our contribution is the empirical demonstration that this extrapolation is valid, and spurious states indeed exist in practically used settings.
In the AM literature, specifically on Hopfield networks, the concept of spurious states is often entangled with the phenomenon of mixture states, which are linear combinations of a certain number of memory vectors. While in certain kinds of AM models these two notions can be closely related, it is important to emphasize that they are different concepts. Spurious states are defined by the existence of a basin of attraction around them. They may or may not be mixture states. The main message of our work is to establish that spurious states exist in DMs. We do not claim that these spurious states are necessarily mixture states. It is up to the model, defined by the score modeling neural network, how the spurious states are constructed from the training data.
Careful characterization of the memorization-generalization transition and the computational aspects of spurious states that define the onset of this transition may offer future insights for guiding the development of methods that can mitigate DMs’ memorization.
4.1  Memorization-to-Generalization Transition
@@
Toy Model. An instructive example, which can reinforce the energy connection between Eqs. (5) and (6), can be shown via a simple 2-dimensional toy model exhibiting many aspects of the memorization-to-generalization transition. Imagine that the training data lies on a unit circle, see Fig. (2). We are interested in exploring how the shape of $E^{\text{DM}}$ (6) changes as the number of training data points increases. Consider the case when the number of training data points is infinite or $K\rightarrow\infty$. Using Eq. (6), the energy of this toy model can be derived as
$\begin{split}E^{\text{AM}}(R,\phi)&=R^{2}+1-\frac{1}{\beta}\log\big[I_{0}(2\beta R)\big]\underset{\beta\rightarrow\infty}{\approx}(R-1)^{2},\end{split}$ (8)
where $R$ is the radius of the unit circle, $\phi$ is the polar angle, and $I_{0}(\cdot)$ is a modified Bessel function of the first kind. The dependence on $\phi$ in Eq. (15) completely disappears from the final result, since the local minima of the resulting energy function form a continuous manifold, described by a parabola centered around $R=1$ as the inverse temperature $\beta\rightarrow\infty$. This behavior describes the fully generalized phase of this toy model. Please refer to the Supplementary for Eq. (15) derivation.
To test the resemblance of $E^{\text{DM}}$ and $E^{\text{AM}}$, we trained a set of DMs in this precise setting. Each model is trained on a dataset of size $K$. The resulting energy landscapes are shown in Fig. (2). For a small data size of $K=2$, the DM exhibits memorization. The local minima of the energy correspond to the training data points. Importantly, at $K=9$ we can observe the first signs of spurious states. At this stage, the model begins to learn emergent (different from the training data) local minima of the energy. Subsequent increase of the size of the training set leads to fully generalized behavior, illustrated at $K=1000$. At this stage, all of the samples from the model live in close proximity of the exact data manifold. The right panel shows the analytical expression of the energy landscape defined by Eq. (15). Thus, the conventional diffusion modeling pipeline agrees very well with the theoretical prediction of the empirical energy.
Figure 3: Different sample types across the memorization-to-generalization transition for CIFAR10. The grey histogram shows the distances between synthetic samples and their nearest neighbors from the synthetic set $\mathsf{S}^{\prime}$. The threshold $\delta_{s}$ defines a boundary between the two peaks. The olive histogram depicts the distances from the synthetic samples to their closest neighbor from the training set $\mathsf{S}$, with threshold $\delta_{m}$ separating the two peaks. Memorized samples are located in the left peak of the olive histogram, below $\delta_{m}$. In contrast, generalized and spurious samples appear to the right of $\delta_{m}$ in the olive histogram. Examples of the generated samples forming each of the four peaks of the histograms are shown in the inset frames. For each generated sample, top-4 nearest neighbors from the training set are shown in the top row, and those from the synthetic set are shown in the bottom row. Training set size $K=7310$ is used in this figure, but the discussed phenomena are general and largely independent of this specific value. The fraction of the memorized, spurious, and generalized samples in the pool of all generated samples is shown in the bottom left panel as a function of the training set size. The inset shows amplified spurious fraction (green curve).
Natural Datasets. However, the main goal of our work is to also establish the existence of spurious states in DMs trained on natural high-dimensional datasets. With this goal in mind, we design detection metrics that can classify any generated sample into one of three categories – memorized, spurious, and generalized samples – to aid in our analyses. Fundamentally, these metrics defined below rely on two datasets: the training dataset $\mathsf{S}$ used to train the DM, and the synthetic dataset $\mathsf{S}^{\prime}$ generated from the fully trained DM. The size of $\mathsf{S}^{\prime}$ is assumed to be much bigger than $\mathsf{S}$.
Our core intuition is rooted in the energy landscape, where its topography dictates the nature of the generated samples, as shown in Fig. (1A). Each sample type corresponds to a distinct feature of this landscape. Memorized samples originate from the basins of attraction the model creates around individual data points from the training set $\mathsf{S}$. Because these are strong (having large basins), the generation process frequently converges to exact memorized samples at small data sizes, producing outputs in the synthetic set $\mathsf{S}^{\prime}$ that are perfect duplicates of training data. In contrast, spurious samples arise from new emergent basins of attraction, formed during training, which do not correspond to any specific sample in $\mathsf{S}$. These emergent minima are also stable attractors, which explains why they appear as duplicates within $\mathsf{S}^{\prime}$ but are absent from $\mathsf{S}$. Finally, generalized samples are drawn from flatter regions of the energy landscape. Since these areas support a continuum of low-energy states rather than a single sharp minimum, the probability of landing on the exact same point twice is negligible, resulting in unique novel creations.
With this intuition in mind we examine two histograms, see Fig. (3). First, for every element of $\mathsf{S}^{\prime}$, their distance to their closest nearest neighbor from $\mathsf{S}$ is shown by the olive histogram. Second, for every element of $\mathsf{S^{\prime}}$, their distance to their closest nearest neighbor in $\mathsf{S}^{\prime}$ (excluding itself) is shown by the gray histogram. Both histograms exhibit a clear bimodal shape, indicating that at least two groups of samples are present in $\mathsf{S}^{\prime}$. This bimodality makes it possible to select the memorized $\delta_{m}$ and spurious $\delta_{s}$ thresholds to separate the samples forming the two peaks in each histogram. The left peak of the olive histogram identifies memorized samples by their small distance to the training set $\mathsf{S}$, leaving spurious and generalized samples in the right peak. The gray histogram then distinguishes truly novel patterns: its left peak contains high-frequency duplicates (both memorized and spurious), while its right peak consists of unique generalized samples. This bimodality provides an empirical basis for defining our detection metrics below.
Following Yoon et al. (2023), we define the memorization detection metric $\mathcal{M}$. A sample $\hat{{\mathbf{x}}}\in\mathsf{S}^{\prime}$ is considered memorized if its distance to its nearest neighbor ${\mathbf{x}}_{1}$, from the training set $\mathsf{S}$, falls below a threshold $\delta_{m}\in\mathbb{R}$:
$\mathcal{M}(\hat{{\mathbf{x}}},\mathsf{S})=\mathbb{I}\bigg(d(\hat{{\mathbf{x}}},{\mathbf{x}}_{1})\leq\delta_{m}\bigg),$ (9)
where $\mathbb{I}$ represents the indicator function and $\delta_{m}$ is a threshold derived from the olive distance histograms, see Fig. (3). Second, a sample $\hat{{\mathbf{x}}}\in\mathsf{S}^{\prime}$ is spurious if it is not a memorized state but is a duplicate of the synthetic set $\mathsf{S}^{\prime}$. This is detected by the metric $\mathcal{S}$, which identifies non-memorized samples that are close to their respective nearest neighbor ${\mathbf{x}}^{\prime}_{1}$ within the synthetic set $\mathsf{S}^{\prime}$:
$\mathcal{S}(\hat{{\mathbf{x}}},\mathsf{S},\mathsf{S}^{\prime})=\mathbb{I}\bigg(d(\hat{{\mathbf{x}}},{\mathbf{x}}^{\prime}_{1})\leq\delta_{s}\bigg)\land\neg\mathcal{M}(\hat{{\mathbf{x}}},\mathsf{S}),$ (10)
where $\delta_{s}\in\mathbb{R}$ is a threshold value chosen based on the gray distance histogram. Finally, a sample $\hat{{\mathbf{x}}}$ is generalized if it is neither memorized nor spurious:
$\mathcal{G}(\hat{{\mathbf{x}}},\mathsf{S},\mathsf{S}^{\prime})=\neg\mathcal{M}(\hat{{\mathbf{x}}},\mathsf{S})\land\neg\mathcal{S}(\hat{{\mathbf{x}}},\mathsf{S},\mathsf{S}^{\prime}).$ (11)
Figure 4: Fractions of memorized, spurious, and generalized samples in synthetic sets across training sizes and datasets. As the training data size $K$ increases, memorization decreases and the fraction of generalized samples steadily increases, see top row of Panel (A). The fraction of spurious patterns rises and decreases at the boundary between the memorization and generalization phases, see bottom row of Panel (A). Panel (B) illustrates the average log-volume of the basins of attraction for memorized and spurious samples being statistically larger than that of generalized samples across all datasets. The shaded regions indicate standard deviation of the log-volume.
[…]
Furthermore, motivated by Biroli et al. (2024); Raya and Ambrogioni (2024); Li and Chen (2024), for these three distinct states we measure the critical time $t_{c}$, where the reverse process (2) can still recover a target pattern $\hat{{\mathbf{x}}}$ from its perturbation ${\mathbf{x}}_{t_{c}}$ within an error margin. The time $t_{c}$ helps define the radius $R(\hat{{\mathbf{x}}},{\mathbf{x}}_{t_{c}})=\lVert\hat{{\mathbf{x}}}-{\mathbf{x}}_{t_{c}}\rVert$ of the basin of the attractions of a sample $\hat{{\mathbf{x}}}$, which is the Euclidean norm of the difference between $\hat{{\mathbf{x}}}$ and its recoverable perturbation ${\mathbf{x}}_{t_{c}}$. The volume of this basin is simply defined by the $N$-ball volume:
$V(\hat{{\mathbf{x}}},{\mathbf{x}}_{t_{c}})=\frac{\pi^{\frac{N}{2}}}{\Gamma(\frac{N}{2}+1)}R(\hat{{\mathbf{x}}},{\mathbf{x}}_{t_{c}})^{N},$ (12)
where $\Gamma(\cdot)$ is the Gamma function. For numerical stability, we show the results of the average logarithmic of Eq. (34) in Fig. (4B) computed from samples belonging to each distinct state.
Confirming the phenomenon observed in our toy model and core intuition, our experiments on these datasets demonstrate a clear three-phase transition from memorization to generalization. As shown in Fig. (4A), models trained in the small data regime predominantly replicate training samples. As the training data size $K$ surpasses a critical capacity, memorization sharply declines, and spurious samples emerge, signaling the onset of generalization. Further increase of $K$ diminishes the fraction of both memorized and spurious samples, and the model enters the full generalization regime, frequently producing unique novel patterns, see Fig. (1B) for the visualization of these three distinct patterns.
Additionally, our results reveal a distinct hierarchy in the basins of attraction for the three sample types. Memorized samples consistently exhibit the largest log-volume, followed by spurious samples, while generalized samples have a much smaller volume, consistent with theoretical expectations. Please refer to Supplementary for more details of this section and additional results.
Figure 5: Panel (A) depicts the average singular values of the energy curvature for different sample types and datasets and the shaded region is the standard deviation of the singular values (a particular training set size $K$ is shown). Memorized and spurious samples generally exhibit higher energy curvature, characterized by fewer near-zero singular values than generalized samples. Bottom row of (A) illustrates the average spectra computed for the training data points, where the changing color scheme denotes small to large data sizes, suggesting a drop in curvature (indicated by the decrease in singular values) as $K$ increases. Panel (B) shows candidate examples of memorized, generalized, and spurious samples from a Stable Diffusion model Rombach et al. (2022) trained on the LAION dataset Schuhmann et al. (2022), each corresponding to a distinct curvature signature. The candidate memorized sample has much larger singular values, while the candidate generalized sample has much smaller ones. The selected possible spurious samples have larger singular values than the candidate generalized sample, but smaller than the candidate memorized sample. The common trait between candidate memorized and spurious samples is that both are stable attractors, as demonstrated by their repeated similar generations given different initial points (or noise vectors) and the conditioning on text-prompt. The y-axis is clipped at the value of 1500 to better contrast the shown examples’ spectra.
4.2  Energy Curvature
While the changing fractions of memorized, spurious, and generalized samples constitute a good signature of the memorization-to-generalization transition, the underlying mechanism is also rooted in the geometry of the model’s learned energy landscape, as shown in Figs. (1A) and (2). To elucidate this mechanism, we quantify the energy curvature at the minima corresponding to each sample type obtained at a small time $t_{0}$.
To measure this curvature, we rely on the relationship $\nabla_{{\mathbf{x}}_{t}}\log p_{t}({\mathbf{x}}_{t})=-\nabla_{{\mathbf{x}}_{t}}E_{t}({\mathbf{x}}_{t})$ to analyze the Hessian of the energy function or the Jacobian of the score, evaluated at the location of each generated sample $\hat{{\mathbf{x}}}\in\mathbb{R}^{N}$. Specifically, following Stanczuk et al. (2022); Achilli et al. (2024); Ventura et al. (2024), we numerically approximate the spectra of the energy Hessian by exploiting the tangent space of the model’s learned manifold at a given pattern $\hat{{\mathbf{x}}}$. First, we generate a set of $4N$ perturbations around $\hat{{\mathbf{x}}}$ using a small time $t_{0}$ in Eq. (1). The learned score $s_{\theta}(\hat{{\mathbf{x}}}_{t_{0}},t_{0})$ is then evaluated at each perturbed point $\hat{{\mathbf{x}}}_{t_{0}}$ and the resulting $4N$ score vectors form a matrix $\mathsf{M}\in\mathbb{R}^{N\times 4N}$. Singular value decomposition is then performed on $\mathsf{M}$ to extract its singular values, serving as our measure of the local curvature of $\hat{{\mathbf{x}}}$ in all directions. The curvature spectra for the three sample types for different datasets are illustrated in the top row of Fig. (5A). Please refer to the Supplementary for further experimental details.
While a consistent separation in the curvature spectra of memorized, spurious, and generalized samples is not readily apparent for all training data sizes $K$, likely due to the geometric complexity of the transitional phases, we are still able to observe very well the separation of the curvature spectra between memorized and generalized samples at many sizes of the training dataset (prominently at or after the peak of spurious states). In addition, we repeated a similar analysis for training data points. There is a consistent decrease in the singular values measured at the locations of the training data points as their number $K$ increases, see bottom row of Fig. (5A). This provides a direct geometric evidence of the mechanism behind generalization: the model transitions from creating sharp, high-curvature minima for individual data points in the memorization regime to learning a smoother, lower-curvature manifold that encompasses the entire dataset in the generalization regime.
To validate these geometric findings on a large-scale model, we performed a case study using Stable Diffusion Rombach et al. (2022), trained on the LAION dataset Schuhmann et al. (2022), shown in Fig. (5B). The analysis of handpicked candidate examples of memorized, spurious, and generalized samples from LAION reveals a similar hierarchy of energy curvatures. As expected, a heavily memorized sample occupies a high-curvature minimum, evidenced by large singular values. In contrast, a more generalized sample exhibits a significantly lower-curvature signature, while spurious samples occupy an intermediate geometric space. Notably, the spectra for all generated samples remain relatively sharp, with few near-zero singular values, as similarly found in Jeon et al. (2024). We argue this is a consequence of text conditioning, which constrains the generative manifold, making Stable Diffusion more prone to operate in high-curvature modes, increasing the risk of memorization.
4.3  Relative Energy Across the Transition
Figure 6: Average relative energy gap of each sample type and the training samples in CIFAR10 Krizhevsky et al. (2014) as training data size $K$ grows. The gaps are measured relative to training samples, whose energy is set to zero (energy is defined up to an additive constant). Shaded regions show standard deviation of the gap values.
Alongside our energy curvature analyses, we study how the energy of these three distinct sample types changes as the training data size increases. Following the efficient relative energy computation from Raya and Ambrogioni (2024), we measure the relative energy gap of the three sample types with respect to that of the training data samples. The results are recorded in Figure˜6. Please refer to Appendix˜E for full details on the experiment and computation of the relative energy gap for different sample types, alongside additional results.
For small training data sizes, the energy of the memorized samples is low and of the generalized samples is high. As $K$ increases, the energy of the memorized samples increases, while the energy of the generalized samples decreases. The energy of the spurious states is the lowest among these three distinct sample types during the memorization-generalization transition.
5  Conclusion
In conclusion, we established a novel connection between DMs and DenseAMs by identifying the emergence of creativity (generalization) in DMs with the failure of successful memory recalls via an energy-based associative memory. Using this theory, we empirically identified a novel intermediate phase, spurious states, previously overlooked in related DM literature, which marks the onset of generalization in these models as the training data size increases. Additional analyses of the volume of the basins of attraction on samples generated from DMs and various aspects about their energy help validate our hypothesis delineated in Fig. (1). Our findings on Stable Diffusion suggest that these novel, stable attractors persist even in large models trained on massive datasets. Thus, our framework provides a new tool for understanding the computational mechanisms of DMs, opening new avenues for characterizing their behaviors.
Acknowledgments
The results presented here were obtained while Dmitry Krotov was employed by IBM Research. At the time of the submission Dmitry Krotov is no longer employed by IBM Research. Matteo Negri acknowledges the support of PNRR MUR project PE0000013-FAIR. Bao Pham and Mohammed Zaki acknowledge the support of RPI-IBM Future of Computing Research Collaboration program (FCRC).
Appendix
Appendix A Toy Example
A.1  Overview
@@
For the 2D toy example, our training data points lie on a unit circle. In the case of infinite training data size or $K\rightarrow\infty$, the distribution of these data points can be described by a continuous density of states:
$p({\mathbf{y}})=\frac{1}{\pi}\delta\big(y_{1}^{2}+y_{2}^{2}-1\big).$ (13)
@@
$p({\mathbf{x}})\sim\int\limits_{-\infty}^{+\infty}dy_{1}dy_{2}\ p({\mathbf{y}})\ e^{-\beta\lVert{\mathbf{x}}-{\mathbf{y}}\rVert^{2}_{2}}=e^{-\beta(R^{2}+1)}I_{0}(2\beta R),$ (14)
@@
where $I_{0}(\cdot)$ is a modified Bessel function of the first kind. Thus, the energy of the model is given by
@@
$\begin{split}E^{\text{AM}}(R,\phi)&=R^{2}+1-\frac{1}{\beta}\log\big[I_{0}(2\beta R)\big]\underset{\beta\rightarrow\infty}{\approx}(R-1)^{2},\end{split}$ (15)